Diffusion Morning Report 09/19/07

Ted Berg tberg at stanford.edu
Wed Sep 19 06:35:10 PDT 2007

Good morning All,
    Looking  OK this morning. Tylan 1 is back together and need s test. 
POCl3 and BBr still under observation. GeH4 changed for SiGe. Thermco 
Poly 2 testing. Ebeam down for beam issues.  No other issues at this time.

  	Date: 	9/19/07 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	down  	Tube replaced needs test 	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	yellow 	observation 	  	  	 
Tylan 6-POCl3 	yellow 	smoking 	  	  	 
Tylan 7-TEOS2 	yellow 	  	2.86/30 	  	 
Tylan 8-SiGe 	up 	GeH4 changed 	22.6/30 	  	 
Tylan 9-Poly 	up 	  	24.72/30 	  	 
Tylan 10-Nitride 	up 	update log 	.67/30 	  	 
Tylan 11-BPSG-LTO 	up 	  	12.84/30 	  	 
Tylan 12-FGA 	up 	tube changed 	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	future 	testing for qual 	
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	
Burn Box 	Up 	  	   	  	 
Ebeam 	down  	beam issues 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	A on line 510 psi 	 changeover working well  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20070919/d6e17c23/attachment.html>

More information about the diffusion mailing list